Mr.Euveja has a total amount of 25,000.He invested this in 2 funds {(Bank A paying 6%)&(Bank B paying 8%) both as annual interest. The combined annual interest is 1,800.
HOW MUCH OF THE 25,000 IS INVESTED INEACH FUND???

Answer by mananth(16946)

Investment A 6.00% per annum ---x
Investment B 8.00% per annum ---y

x + y= 25000 ------------------------1
6.00% x + 8.00% y= = $1,800.00
Multiply by 100
6 x + 8 y= = $180,000.00 --------2

Multiply (1) by -6
we get

-6 x -6 y= = -150000.00

Add this to (2)

0 x 2 y= = $30,000.00

divide by 2

y = $15,000.00 investment at 8.00%
Balance $10,000.00 investment at 6.00%
CHECK
$10,000.00 --------- 6.00% ------- $600.00
$15,000.00 ------- 8.00% ------- $1,200.00
Total -------- -------------- $1,800.00
m.ananth@hotmail.ca